Shift report 20011007 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Oct 07 2001 - 16:39:33 EDT

  • Next message: Zbigniew Majka: "Unsual Shift."

    ------------------------------------------------------------
    Date: Sun Oct  7 16:39:32 2001, Shift: 08:00-16:00
    Supervisor: fv, with: 
    Log book#9 pages 111 - 114
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine did work on betasqueeze. two succesful ramps.
       No data taking to HPSS.
    2. Checked effect of trigger 5 (vertex cut). Seems efficient > 99%
       for +-20 cm. 
       Modified SumamryPlot.C to use pdq - can now print from 
       pii5.
    3. Misc problems
       Had to cycle power on rack 4 crate 3- lost triggers.
       The UPS for OPUS indicates errors - 
    	 - Bob&Betty please check 
        MTPC2 still see problems.
    
    stared run 5221 at beginnign of necxt shift; 
       12 deg / 3 deg T3 turned down contributed ~30% 
       of datavolume t4,t5 may be useful for calibration & check
       out.
    
    8.00  Status.
          MRS at 45 deg. Have .3M at 4A continue.
          FS at  12/3 deg. Need another ~1.7 for 1/5A. 
          Then move 
    
    9.45 (FV) Modified the SummaryPlots.C to use the ~daq/bin/pdq 
         which seems to workjust fine. So brntuple can be run from pii5.
         The printer lpd-deamon problem will still have to be looed at, again.
    
    10.15 (FV) Machine will attempt beta*=2 stores, and have been trying 
          since ~8.50. 
          An apparent succesfull ramp is underway
    
    10.20 (FV) Store with beta(*)=2 but no 200 MHz. You gain some, but loose 
          at the same time Hopefully they will get this right later.
          Clearly steering is needed too.
          Our ZDC rates are 22 Phenix 540, star+phobos==0
    
    10.33 Since the 200MHz is not on, the rates could well be high, and the
          vertex trigger really seem fully efficient(>99%)  for T(BB)<1.5 nsec and
          for BB(Multsum) > 20- I trust it will be most useful to turn this one
          at least for this store.
    
    10.40 Machine folks are working on tunes (Blue in particular) to improve 
          lifetime before steering. 
    
    11.00 This store seem to be a dedicated machine development study - this is 
          good for us if the beta*=2 gets well developed.
          records some runs to spool only for effeciency studies of trigger.
          The vertex is way off even though we coincidenetly now have a coll rate
          of ~ 150-> 100. See attached plot of BB vertex (in cm).
          (peak at ~-40 cm. ~1/4 of rate within +-30 cm.
    
    11.06 Beam dumped --.
          --later message It was actully lost.
    
    
    11.44 Updated some html links in op pages that were not correct.
          Added .htaccess to the daq public_html pages. The stuff on the DAQ
          pages is not public, but for collaborators only. E.g. some pages 
          on the daq page gives controls to online and control stuff, though 
          still through anotherlayer of password to power in IR.
    
    11.49 Another RHIC ramp in progress.
    
    12.26 The store is being used for machine developments.
          Will look at a few things in DAQ.
       
    13.20 Had no trigger 4. It turned out the camac crate 3 in rack 4 that 
          distributes trigger signals were stuck. Recycling the power cleared
          things.
          The vertex in this store in now shiftewd to the right (+50 cm).The rates
          are also six times smnaller than the Phenix values.
    14:00 RHIC is finding a better tune/steering for us. The rates are up to ~247 
          while Phenix and Star is about 400. The vertex is also much better 
          centered.
    14:18 For on-going test run - with data not being recorded -mtpc2 still 
          exhibits error. This should be checked during the maintnance period
          with pulsers to see if we can identify the problem.
    
    14:20 Record to spool run#5217 with tr(1) = 1; tr(3)=0 ; tr(4)=20 and tr(5)= 1
          TPCs on. DC's off.
          ZDC ~ 200  BI ~17 Y 18
          Rhic is steering at Phobos.
    
          Daq rates ~43/sec  24% dead.
    
          Took about 63 K events.
    
    15:05 stooped this run
    
    15:30 The RHIC stored was lost/dumped.
          Brahms rates still ~50% other even though centered well enough.
          
    15:45 Learned from rhic e-log that cause of dump is to refill for experiment
          store. The claim is also the the Brahms rates are lower due to a bigger
          effect of crossing angle when beta-star is smaller.
    
    15:50 #Beam hours this shift==0 (beam studies).
    
    
    Detailed e-log
    
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Oct 07 2001 - 16:40:20 EDT